November 1997  ICTCM X (Chicago)  33

Ongoing Efforts

• Implementation

– continued development of specialized and general CAS tools for symbolic, numeric, and graphic analysis of BVPs

– collect realistic and interesting examples of BVPs for use as homework, in-class examples, and projects

– generalize and extend BVPsolve and fBVPsolve (and shoot)

• General Goals – BVPs deserve more attention than they currently receive – should some material be omitted from traditional courses to

allow discussion of BVPs?

– how do BVPs fit into newer courses that emphasize dynamical systems?